You are here

Beam delivery update: Last update of 2014?

header

12 December 2014

Not much has changed since last night – the cyclotron is still delivering close to 100 µA to BL2C4 for isotope production and protons as needed to BL1B for PIF. BL2A is still off while the next experiment is being set up; the plan is for ISAC Ops to request proton beam and to establish the tune through the polarizer over the weekend in advance of RIB delivery on Monday. That experiment will run until ~18:30 on December 21, at which time ISAC Ops will take the beam back to carry out target scans and the end-of-target checklist before shutting down for the year. 

Please note that next week is the last week of operation for 2014. As I understand it, isotope production at STF will end Tuesday morning when PIF resumes on BL2C1; there is also beam development planned for Thursday’s day shift. Otherwise, cyclotron operation will continue until the start of shutdown at 20:30 on Sunday, December 21.

On a related note, this will likely be the last beam delivery update of the year as I will be on vacation starting Monday (though I may still be around early in the week). There may be occasional updates during shutdown, but regular updates won’t resume until  we start getting ready to resume operation in the spring. Happy holidays, and all the best in the new year.

-- ACM